An office building is expected to create operating cash flows of $25,500 a year for three years, based on tenants rental income. The purchase of the fixed assets for this building will cost $53,000. These assets will have no value at the end of the project. Ar additional $4,000 of net working capital will be required throughout the life of the project. Calculate the net present value of this project if the required rate of return is 13 percent?
Answer #1


Net Present Value = $5,981.59

Year

Cash flows = CF

Working capital adjustment = WC

Net cash flow = NCF

Discount factor = Df = 1/(1+Rate)^Year

Present Values

0

-$53,000.00

-$4,000.00

-$57,000.00

                      1.00000

-$57,000.00

Y

CF

WC

NCF = CF + WC

Df = 1/(1+13%)^Y

=Df x NCF

1

$25,500.00

$25,500.00

                      0.88496

$22,566.37

2

$25,500.00

$25,500.00

                      0.78315

$19,970.24

3

$25,500.00

$4,000.00

$29,500.00

                      0.69305

$20,444.98

Total = NPV = PW=

$5,981.59